OpenStudy (anonymous):

Convert 50 years into seconds. Express your answer in scientific notation

OpenStudy (perl):

how many seconds are in a minute how many minutes are in an hour how many hours are in a day how many days are in a years

OpenStudy (perl):

multiply all these

OpenStudy (perl):

that is the number of seconds in a year , and you have 50 years

OpenStudy (aripotta):

\[50yrs * \frac{ 364days }{ 1yr }*\frac{ 24hrs }{ 1day }*\frac{ 60\min }{ 1hr }*\frac{ 60\sec }{ 1\min }\]

OpenStudy (anonymous):

Thanks